# II. Carbon-enhanced metal-poor starsthanks: All M3DIS 3D RHD CEMP models are available under https://nlte.mpia.de

### v1

@@ Appendix A Model parameter selection
Figure 11: 2D histogram of the dataset used in this paper in the $\rm T_{eff}$ vs. $\rm\log g$ plane. The $\rm\log g$-limit of $2.7$ applied in this work is shown as a black, dashed line. The black, solid line represents the mean in bins of constant $\rm\log g$. Red and blue crosses are placed at the location of the 3D M3DIS models. The insets show the temperature distribution at the optical surface of the respective model with $\rm[Fe/H]=-5$. To show the different physical scales involved, a blue square is added to the inset of the sub-giant, showing the size of the main-sequence model in comparison. To estimate the dependence of 3D $\rm A(C)$ corrections on effective temperature, additional models are created at the locations highlighted with green crosses. See text.
@@
To justify the selection of model parameters we show the $\rm T_{eff}$-$\rm\log g$ distribution of the stars in our dataset in Fig. 11. In this figure, red and blue crosses denote the positions of the sub-giant and MS models, respectively. Given the numerical complexity of the analysis in the paper, we restricted the number of parameters and constructed 3D M3DIS model atmospheres for these two $\rm T_{eff}$-$\rm\log g$ combinations only (see Tab. 2), which are chosen to best represent the stellar sample. To quantify the influence of this selection on the $\rm 3D-1D$ abundance corrections, we additionally compute a range of 3D M3DIS models in-between, with a single representative chemical composition of $\rm[Fe/H]=-5,[C/Fe]=0$ and $\rm\log g=4.0$ (green crosses). We create these models for a range of different effective temperatures between $\rm 5000\,K$ and $\rm 6250\,K$ to estimate the effect of $\rm T_{eff}$ on abundance corrections, while keeping $\rm\log g$ fixed. The results are shown in the bottom panel of Fig. 12.
@@
Figure 12: Top: Formation height of a representative CH line in the atmospheres of 3D M3DIS models at $\rm[Fe/H]=-5,\ \log g=4.0$. Bottom: 3D scaled-solar – 1D $\rm A(C)$ abundance corrections as a function of $\rm T_{eff}$, derived from spectra using the same atmosphere models as in the top panel and applying $\rm[C/Fe]=3.0$ as reference abundance, as in Fig. 6
@@
When comparing with Fig. 6, it may be noted that the $\rm 3D\ \text{scaled-solar}-1D$ $\rm A(C)$ abundance corrections overall are very similar to the corrections computed for MS and sub-giant models, indicating that there is no significant effect on the overall distribution of CEMP stars that influences the conclusions drawn in this paper. At the same $\rm T_{eff}=5750\,K$, the difference between the MS ($\rm\log g=4.5$) and the additional model ($\rm\log g=4.0$) is negligible. Corrections for the sub-giant are slightly larger than the results at $\rm T_{eff}=5250\,K$, which is expected and may be attributed to its significantly lower surface gravity. Stars near $\rm\log g=4.0$ align more closely with the MS model than with the sub-giant model regarding stellar parameters, so this discrepancy is not relevant for the final distribution.
@@
Investigating the corrections as a function of effective temperature, it becomes apparent that $\rm\Delta A(C)$ increases (in absolute values) as $\rm T_{eff}$ increases. Considering that the strength of molecular spectral lines weakens with increasing temperature, this may seem counter-intuitive at first. To understand this behaviour qualitatively, we additionally include the formation height of a representative CH line in the G-band for different $\rm T_{eff}$ in the upper panel of Fig. 12, in 1D (red, dashed) as well as in 3D (blue, solid). Investigating the 1D trend it may be noted that – as expected – the formation height of CH lines decreases, i.e. shifts towards the stellar interior as the atmosphere gets hotter. However, in 3D the opposite is the case; here the negative temperature gradient in the upper layers of the atmosphere – which is not present in 1D models in radiative equilibrium – still provides conditions favourable for molecule formation in cooler layers further out, even though the overall temperature at the optical surface increases. The formation height hence shifts in the opposite direction, which causes the structural and spectral differences between 3D and 1D to increase. Because the hottest model we include in the main analysis of the paper has $\rm T_{eff}=5750\,K$, we find that the $\rm A(C)$ correction for stars in our sample around $\rm\log g\sim 4.0,\ [Fe/H]\sim-5$ at the location of the mean ($\rm T_{eff}\sim 6000\,K$) may be underestimated by $\sim\rm 0.2\,dex$. We thus expect that 3D effects become even more significant when the sampling of the parameter space is increased, which further supports the conclusions drawn in this study and underlines the need for 3D model atmospheres for the analysis of metal-poor stars. It remains to be investigated how this temperature dependence behaves as a function of metallicity and with C-enhancement included in the computation of the atmosphere itself.
@@ Appendix B Updated radiative transfer solver
Figure 13: Comparison between the horizontally averaged radiative heating $\rm Q_{r}$, computed at different resolutions for the same model atmosphere using the old, classical RT solver (top) and new, conservative RT solver on a staggered grid (bottom). The new solver reproduces the cooling peak very well already at intermediate resolution, while the classical solver shows the typical undersampling effect, which leads to artificial heating of the model.
We have found that the cost of computing accurate vertical temperature profiles at the surface can be significantly reduced, relative to Eitner et al. (2024), by modifying the centering of RT solutions in the short-characteristics solver. Instead of computing the radiation intensities $\rm I_{\Omega,\nu}$ and source functions $\rm S_{\nu}$ at cell centers, and computing the net radiative cooling from
$Q_{\rm r}=\iint{\rm d}\Omega\,{\rm d}\nu\,\rho\kappa_{\nu}\,(S_{\rm\nu}-I_{\rm\Omega,\nu})\,$, (2)
one can instead compute the radiative intensities $\rm I_{\Omega\nu}$ at cell interfaces, and from these compute the radiative fluxes $F_{i,\nu}$ in each direction $i$, which allows the cooling to be obtained from
$Q_{\rm r}=\int{\rm d}\nu\ \nabla\cdot\mathbf{F_{\rm\nu}}\,$, (3)
where the flux divergence is computed numerically as the differences between incoming and outgoing fluxes for each cell. This method is by construction energy conserving, in the sense that the cooling corresponds exactly to the energy increase of the radiation field leaving the cells, relative to the energy that was entering. Thus, while the previous method ‘samples’ the cooling point-wise, the improved method computes the actual energy loss across cells.

### v2

@@ Appendix A Model parameter selection
Figure 11: 2D histogram of the dataset used in this paper in the $\rm T_{eff}$ vs $\rm\log g$ plane. The $\rm\log g$-limit of $2.7$ applied in this work is shown as a dashed black line. The solid black line represents the mean in bins of constant $\rm\log g$. Red and blue crosses mark the location of the 3D M3DIS models. The insets show the temperature distribution at the optical surface of the respective model with $\rm[Fe/H]=-5$. To highlight the different physical scales involved, the blue square in the inset shows the size of the MS model. To estimate the dependence of 3D $\rm A(C)$ corrections on effective temperature, additional models are created at the locations highlighted with green crosses.
@@
To justify the selection of model parameters, we show the $\rm T_{eff}$-$\rm\log g$ distribution of the stars in our dataset in Fig. 11. Given the numerical complexity of the analysis in the paper, we restricted the number of parameters and constructed 3D M3DIS model atmospheres for these two $\rm T_{eff}$-$\rm\log g$ combinations only (see Table 2), which are chosen to best represent the stellar sample. To quantify the influence of this selection on the $\rm 3D-1D$ abundance corrections, we additionally computed a range of 3D M3DIS models in between, with a single representative chemical composition of $\rm[Fe/H]=-5,[C/Fe]=0$ and $\rm\log g=4.0$ (green crosses). We created these models for a range of different effective temperatures between $\rm 5000\,K$ and $\rm 6250\,K$ to estimate the effect of $\rm T_{eff}$ on abundance corrections, while keeping $\rm\log g$ fixed. The results are shown in the bottom panel of Fig. 12.
@@
Figure 12: Top: Formation height of a representative CH line in the atmospheres of 3D M3DIS models at $\rm[Fe/H]=-5,\ \log g=4.0$. Bottom: 3D scaled-solar – 1D $\rm A(C)$ abundance corrections as a function of $\rm T_{eff}$, derived from spectra using the same atmosphere models as in the top panel and applying $\rm[C/Fe]=3.0$ as the reference abundance, as in Fig. 6
@@
When comparing this panel with Fig. 6, we see that the $\rm 3D\ \text{scaled-solar}-1D$ $\rm A(C)$ abundance corrections overall are very similar to the corrections computed for MS and sub-giant models, indicating that there is no significant effect on the overall distribution of CEMP stars that influences the conclusions drawn in this paper. At the same $\rm T_{eff}=5750\,K$, the difference between the MS ($\rm\log g=4.5$) and the additional model ($\rm\log g=4.0$) is negligible. Corrections for the sub-giant are slightly larger than the results at $\rm T_{eff}=5250\,K$, which is expected and can be attributed to its significantly lower surface gravity. Stars near $\rm\log g=4.0$ align more closely with the MS model than with the sub-giant model regarding stellar parameters, so this discrepancy is not relevant for the final distribution.
@@
Investigating the corrections as a function of effective temperature, it becomes apparent that $\rm\Delta A(C)$ increases (in absolute values) as $\rm T_{eff}$ increases. Considering that the strength of molecular spectral lines weakens with increasing temperature, this may seem counter-intuitive at first. To understand this behaviour qualitatively, we additionally include the formation height of a representative CH line in the G band for different $\rm T_{eff}$ in the upper panel of Fig. 12, in 1D (dashed red) as well as in 3D (solid blue). Investigating the 1D trend, we see that – as expected – the formation height of CH lines decreases, i.e. shifts towards the stellar interior, as the atmosphere gets hotter. However, in 3D the opposite is the case; here the negative temperature gradient in the upper layers of the atmosphere – which is not present in 1D models in radiative equilibrium – still provides conditions favourable for molecule formation in cooler layers further out, even though the overall temperature at the optical surface increases. The formation height hence shifts in the opposite direction, which causes the structural and spectral differences between 3D and 1D to increase. Because the hottest model we included in the main analysis has $\rm T_{eff}=5750\,K$, we find that the $\rm A(C)$ correction for stars in our sample around $\rm\log g\sim 4.0,\ [Fe/H]\sim-5$ at the location of the mean ($\rm T_{eff}\sim 6000\,K$) may be underestimated by $\sim\rm 0.2\,dex$. We thus expect that 3D effects become even more significant when the sampling of the parameter space is increased, which further supports the conclusions drawn in this study and underlines the need for 3D model atmospheres for the analysis of metal-poor stars. It remains to be investigated how this temperature dependence behaves as a function of metallicity and with C-enhancement included in the computation of the atmosphere itself.
@@ Appendix B Updated radiative transfer solver
Figure 13: Comparison of the horizontally averaged radiative heating ($\rm Q_{r}$), computed at different resolutions for the same model atmosphere using the old, classical RT solver (top) and the new, conservative RT solver on a staggered grid (bottom). The new solver already reproduces the cooling peak very well at intermediate resolution, while the classical solver shows the typical under-sampling effect, which leads to an artificial heating of the model.
We have found that the cost of computing accurate vertical temperature profiles at the surface can be significantly reduced, relative to Eitner et al. (2024), by modifying the centring of RT solutions in the short-characteristics solver. Instead of computing the radiation intensities $\rm I_{\Omega,\nu}$ and source functions ($\rm S_{\nu}$) at cell centres, and computing the net radiative heating from
$Q_{\rm r}=\iint{\rm d}\Omega\,{\rm d}\nu\,\rho\kappa_{\nu}\,(I_{\rm\Omega,\nu}-S_{\rm\nu})\,$, (2)
one can instead compute the radiative intensities $\rm I_{\Omega\nu}$ at cell interfaces, and from these compute the radiative fluxes $F_{i,\nu}$ in each direction $i$, which allows the heating to be obtained from
$Q_{\rm r}=-\int{\rm d}\nu\ \nabla\cdot\mathbf{F_{\rm\nu}}\,$, (3)
where the flux divergence is computed numerically as the differences between incoming and outgoing fluxes for each cell. This method is by construction energy-conserving, in the sense that the cooling corresponds exactly to the energy increase in the radiation field leaving the cells, relative to the energy that was entering. Thus, while the previous method ‘samples’ the cooling point-wise, the improved method computes the actual energy loss across cells.